OpenStudy (anonymous):

suppose ln2= r and ln5= s . use properties of logarithms to write the logarithm in terms of r and s. Ln2.5

OpenStudy (anonymous):

use ln2+ln5=ln(2*5)

OpenStudy (anonymous):

i think you should use \[\ln(2.5)=\ln(\frac{5}{2})=\ln(5)-\ln(2)=s-r\]

OpenStudy (anonymous):

so is it s+r or s-r?

OpenStudy (anonymous):

its s-r if 2.5is 5/2 or s+r if 2.5 is 2*5=10

OpenStudy (anonymous):

Ok thanks i think i will write s+r